# An upper bound for the number of smooth values of a polynomial and its applications

### v1

@@ Abstract.
We provide a new upper bound for the number of smooth values of a polynomial with integer coefficients. This improves Timofeev’s previous result unless the polynomial is a product of linear polynomials with integer coefficients. As an application, we fix an error in the proof of a result of Cassels which was used to prove that the Hurwitz zeta-function with algebraic irrational parameter has infinitely many zeros on the domain of convergence. We also apply the main result to a problem on primitive divisors of quadratic polynomials.
@@ 1.1. Statement of the main results
holds in the range $e\leq u\leq(\log{x})(\log\log{x})^{-1}$ for sufficiently large $x$, where $c(f)$ is a positive constant depending on the degree and coefficients of $f(t)$. For more general polynomials, Timofeev [18] obtained that
@@
holds for any $u\geq 1$. This is also useful in applications of Theorem 1.1.
@@ 1.2. Note to “Footnote to a note of Davenport and Heilbronn”
In Section 5, we prove the following result by applying Theorem 1.1.
@@
Theorem 1.2 was essentially obtained by Cassels [2]. He applied this kind of result to study the zeros of the Hurwitz zeta-function $\zeta(s,\alpha)$ for $\operatorname{Re}(s)>1$, where $\alpha$ is an algebraic irrational number with $10<\alpha<$ 1. Indeed, he showed that $\zeta(s,\alpha)$ has infinitely many zeros in the strip $1<\operatorname{Re}(s)<1+\delta$ for any $\delta>0$, which extends the classical result of Davenport and Heilbronn [4]. However, the proof in [2] contains an error which cannot be easily fixed as explained in [16]. See Remark 5.5 for more details of the difference between the methods of [2] and this paper.
Remark that the formulation of Theorem 1.2 is due to Worley [20], and the original statement by Cassels in [2] is slightly different. We discuss how Cassels’ result is proved at the end of Section 5.
@@ 1.3. Primitive divisors of $n^{2}+b$
Another application of Theorem 1.1 is about the primitive divisors of quadratic polynomials. Let $(A_{n})$ be a sequence of integers. We say that an integer $d>1$ is a primitive divisor of $A_{n}$ if $d\mid A_{n}$ and $(d,A_{m})=1$ for all nonzero terms $A_{m}$ with $m<n$. Then we define 111 In the previous papers, $R_{b}(x)$ was denoted by $\rho_{b}(x)$, and usually $\rho_{1}(x)$ was abbreviated to $\rho(x)$. We change the symbols so as to keep that $\rho$ represents the Dickman function.
@@
for $b\in\mathbb{Z}$. Suppose that $-b$ is not an integer square. Then Everest and Harman [6] conjectured that $R_{b}(x)\sim(\log{2})x$ as $x\to\infty$. They also proved
@@
Li [14] also announced an even better upper bound $R_{1}(x)<0.8437x$. In this paper, we provide a new lower bound for $R_{b}(x)$.
@@
Suppose that $-b$ is not an integer square. Then the inequality
@@
The proof of Theorem 1.3 is presented in Section 6. It should be noted that the method for the proof is applicable only to the lower bound for $R_{b}(x)$, and we have no contribution to the upper bound.
For $b=1$, Theorem 1.3 is also related to a problem on arctangent relations which was originally studied by Todd [19]. For $n\in\mathbb{Z}_{\geq 1}$, we say that $\arctan n$ is reducible if there exist integers $f_{1},\ldots,f_{k}$ and $n_{1},\ldots,n_{k}$ with $1\leq n_{j}<n$ such that
@@
It would be worth noting that $N(x)<0.8437x$ also follows from Li’s work [14].
@@ Acknowledgments
The author is especially grateful to Yuta Suzuki for sending a copy and translation of Timofeev’s paper [18]. Many discussions with him were very fruitful. The author would also like to thank Hiroshi Mikawa. Theorems 1.3 and 1.4 were motivated by his lecture at Hachiouji Number Theory Seminar 2024. The work of this paper was supported by JSPS Grant-in-Aid for Early-Career Scientists (Grant Number 24K16906).
@@ 2. Strategy and the initial estimates
where $[k_{1},k_{2}]$ denotes the largest common multiple of $k_{1}$ and $k_{2}$.
@@ 3. An inductive process
(3.2) $S(\kappa)\leq\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\mid f(n)\end{subarray}}1=\sum_{\begin{subarray}{c}u\pmod{\kappa}\\ f(u)\equiv 0\pmod{\kappa}\end{subarray}}\sum_{\begin{subarray}{c}z<n\leq x\\ n\equiv u\pmod{\kappa}\end{subarray}}1\leq\omega_{f}(\kappa)\left(\frac{x-z}{\kappa}+1\right),$
@@
Let $f(t)\in\mathbb{Z}[t]$ be a polynomial such that $f(t)\to\infty$ as $t\to\infty$, and $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 1$. Suppose that $f(z)>x$ is satisfied, and put $h=x-z$. Then we have
@@
Let $f(t)\in\mathbb{Z}[t]$ be a polynomial such that $f(t)\to\infty$ as $t\to\infty$, and $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 1$. Suppose that $f(z)>x$ is satisfied, and put $h=x-z$. Then inequality (2.1) holds with
@@
×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
@@
×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{i+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{i}\leq h,~{}\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{i}}<k_{i+1}\leq f(x)\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{i+1})\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{i+1})$
@@
×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~{}(j\geq 2)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\sum_{\begin{subarray}{c}1\leq\lambda\leq\frac{h}{k_{1}\cdots k_{m}}\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ k_{1}\cdots k_{m}\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
@@
×$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\sum_{\begin{subarray}{c}\frac{h}{k_{1}\cdots k_{m}}<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\omega_{f}(k_{1}\cdots k_{m}\lambda).$
@@
×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}1\leq\lambda\leq\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
@@
×$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\omega_{f}(k_{1}\cdots k_{m}\lambda).$
@@
Note that the function $\omega_{f}(k)$ is multiplicative but is not completely multiplicative in general. Hence we do not expect that $\omega_{f}(k_{1}\cdots k_{m})=\omega_{f}(k_{1})\cdots\omega_{f}(k_{m})$ holds unless $k_{1},\ldots,k_{m}$ are pairwise coprime. 222 It seems to be misunderstood often in [18] that the identity $\omega_{f}(k_{1}\cdots k_{m})=\omega_{f}(k_{1})\cdots\omega_{f}(k_{m})$ holds for any $k_{1},\ldots,k_{m}\in\mathbb{Z}_{\geq 1}$. One can fix the errors with the same argument presented in this paper using Lemma 4.2. Then we use the following lemma which helps us estimate multiple sums involving $\omega_{f}(k_{1}\cdots k_{m})$.
@@
Suppose that $f(t)\in\mathbb{Z}[t]$ is a product of distinct irreducible factors $f_{1}(t),\ldots,f_{g}(t)$ over $\mathbb{Z}[t]$ of degrees $d_{1},\ldots,d_{g}\geq 1$, respectively. Let $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 2$. Suppose further that $f(t)\to\infty$ as $t\to\infty$, and that $f(z)>x$. Put $h=x-z$. Then there exists a constant $A(f)>1$ depending only on $f(t)$ such that, for any $m\in\mathbb{Z}_{\geq 1}$ and any $i=1,\ldots,m$, we have
@@
by recalling that $h=x-z<x$. Applying (4.1) and (4.2) of Lemma 4.1, we have
@@
Suppose that $f(t)\in\mathbb{Z}[t]$ satisfies the same conditions as in Proposition 4.3. Let $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 2$. Put $h=x-z$. Then there exists a constant $A(f)>1$ depending only on $f(t)$ such that, for any $m\in\mathbb{Z}_{\geq 1}$ and any $i=1,\ldots,m$, we have
@@
×$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~{}(j=1,2)\\ P^{+}(k_{j})\leq y~{}(j\geq 3)\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m+1})}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
@@
×$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq y~{}(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
@@
Then we divide the sums into those over $k_{1},\ldots,k_{m+1}\geq 1$ such that $(k_{j},\Delta_{f})=1$ if $j\in S$ and $(k_{j},\Delta_{f})>1$ otherwise, where $S$ is a subset of $\{1,\ldots,m+1\}$. Recall that $[k_{1},k_{2}]k_{3}\cdots k_{m+1}=k_{1}k_{2}k_{3}\cdots k_{m+1}$ holds if $(k_{1},k_{2})=1$. Therefore we obtain
@@
×$\qquad\times\frac{xg^{[u]}}{d(d-1)^{[u]-1}u^{[u]}}\left(1+O\left(\frac{1}{\log\log{x}}\right)\right)$
@@
5. Proof of Theorem 1.2
Let $\alpha$ be an arbitrary algebraic irrational number, and $K=\mathbb{Q}(\alpha)$ the algebraic field generated by $\alpha$ over the rationals. Then we define $P_{K}$ as the set of all prime ideals $\mathfrak{p}$ of $K$ such that the following conditions are satisfied:
@@
In other words, we have $\operatorname{N}(\mathfrak{p})=p$ and $\mathfrak{p}^{2}\nmid(p)$ for a rational prime $p$ if $\mathfrak{p}\in P_{K}$. Let $\mathfrak{a}$ be the ideal denominator of the fractional ideal $(\alpha)$ of $K$. Then we factorize the ideal $(n+\alpha)\mathfrak{a}$ for $n\in\mathbb{Z}$ as $(n+\alpha)\mathfrak{a}=\mathfrak{a}_{n}\mathfrak{b}_{n}$, where
$\mathfrak{a}_{n}=\prod_{\mathfrak{p}\in P_{K}}\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)}\quad\text{and}\quad\mathfrak{b}_{n}=\prod_{\mathfrak{p}\notin P_{K}}\mathfrak{p}^{v_{\mathfrak{p}}(n+\alpha)},$
with $v_{\mathfrak{p}}(n+\alpha)\in\mathbb{Z}_{\geq 0}$ such that $v_{\mathfrak{p}}(n+\alpha)=0$ for all but finitely many of the prime ideals $\mathfrak{p}$. We begin by proving the following preliminary lemmas.

### v2

@@ Abstract.
We prove a new upper bound for the number of smooth values of a polynomial with integer coefficients. This improves Timofeev’s previous result unless the polynomial is a product of linear polynomials with integer coefficients. As an application, we provide another proof for a result of Cassels which was used to prove that the Hurwitz zeta-function with algebraic irrational parameter has infinitely many zeros on the domain of convergence. We also apply the main result to a problem on primitive divisors of quadratic polynomials.
@@ 1.1. Statement of the main results
holds in the range $e\leq u\leq(\log{x})(\log\log{x})^{-1}$ for sufficiently large $x$, where $c(f)$ is a positive constant depending on the degree and coefficients of $f(t)$. See also a recent article of Suzuki [18] for an improvement of Hmyrova’s estimate. For more general polynomials, Timofeev [19] obtained that
@@
holds for any $u\geq 1$. While this may seem like only a slight improvement over the previous result, we will see in Sections 1.2 and 1.3 that it has valuable applications to several issues.
@@ 1.2. Note to “Footnote to a note of Davenport and Heilbronn”
A lower bound for $C_{\alpha}(x)$ was used to study the zeros of the Hurwitz zeta-function $\zeta(s,\alpha)$ for $\operatorname{Re}(s)>1$, where $\alpha$ is an algebraic irrational number with $0<\alpha<1$. Indeed, Cassels [2] applied the estimate $C_{\alpha}(x)>0.51x$ for large $x$ to showed that $\zeta(s,\alpha)$ has infinitely many zeros in the strip $1<\operatorname{Re}(s)<1+\delta$ for any $\delta>0$, which extends the classical result of Davenport and Heilbronn [4]. The relation between $C_{\alpha}(x)$ and the number of smooth values of a polynomial $f(t)\in\mathbb{Z}[t]$ was firstly mentioned by Worley [21], and a more detailed research was presented by Andersson in his unpublished work [1]. In Section 5, we prove the following result.
@@
$C_{\alpha}(x)=x-\Psi_{f}(x,x)+O\left(\frac{x}{\log{x}}\right)$
Therefore, to show a lower (resp. upper) bound for $C_{\alpha}(x)$ is nothing other than to show an lower (resp. upper) bound for $\Psi_{f}(x,x)$. In particular, Theorem 1.1 implies the following result.
Corollary 1.3.
@@
Corollary 1.3 was essentially obtained by Cassels [2] and Worley [21]. Note that the original proofs in [2, 21] should be modified due to an erroneous classification of prime ideals. See [16, Remark 4.10] for the details. The main difference between the methods in [2, 21] and this paper is that the proof of Theorem 1.1 does not require any ideal-theoretic arguments.
@@ 1.3. Primitive divisors of $n^{2}+b$
Another application of Theorem 1.1 is about the primitive divisors of quadratic polynomials. Let $(A_{n})$ be a sequence of integers. We say that an integer $d>1$ is a primitive divisor of $A_{n}$ if $d\mid A_{n}$ and $(d,A_{m})=1$ for all nonzero terms $A_{m}$ with $m<n$. Then we define  [^1] In the previous papers, $R_{b}(x)$ was denoted by $\rho_{b}(x)$, and usually $\rho_{1}(x)$ was abbreviated to $\rho(x)$. We change the symbols so as to keep that $\rho$ represents the Dickman function.
@@
for $b\in\mathbb{Z}$. Suppose that b-b is not an integer square. Then Everest and Harman [6] conjectured that $R_{b}(x)\sim(\log{2})x$ as $x\to\infty$. They also proved
@@
Li [14] also announced an even better upper bound $R_{1}(x)<0.8437x$. In Section 6, we connect $R_{b}(x)$ with the number of smooth values of the quadratic polynomial $f(t)=t^{2}+b$.
@@
Let $f(t)=t^{2}+b$ with b-b not an integer square. Then we have
$R_{b}(x)=x-\Psi_{f}(x,x)+O\left(\frac{x\log\log{x}}{\log{x}}\right)$
Applying Theorem 1.1, we prove a new lower bound for $R_{b}(x)$.
Corollary 1.5.
Suppose that b-b is not an integer square. Then the inequality
@@
For $b=1$, Corollary 1.5 is also related to a problem on arctangent relations which was originally studied by Todd [20]. For $n\in\mathbb{Z}_{\geq 1}$, we say that $\arctan n$ is reducible if there exist integers $f_{1},\ldots,f_{k}$ and $n_{1},\ldots,n_{k}$ with $1\leq n_{j}<n$ such that
@@
It would be worth noting that $N(x)<0.8437x$ also follows from Li’s work [14]. If (1.1) is true for $f(t)=t^{2}+b$ and $u=1$, then Theorem 1.4 yields that
$R_{b}(x)\sim(1-\rho(2))x$
@@ Acknowledgments
The author is especially grateful to Yuta Suzuki for sending a copy and translation of Timofeev’s paper [19]. Many discussions with him were very fruitful. The author would also like to thank Hiroshi Mikawa. Corollaries 1.5 and 1.6 were motivated by his lecture at Hachiouji Number Theory Seminar 2024. The work of this paper was supported by JSPS Grant-in-Aid for Early-Career Scientists (Grant Number 24K16906).
@@ 2. Strategy and the initial estimates
where $k_{1},k_{2}]$ denotes the least common multiple of $k_{1}$ and $k_{2}$.
@@ 3. An inductive process
(3.2) $S(\kappa)\leq\sum_{\begin{subarray}{c}z<n\leq x\\ \kappa\mid f(n)\end{subarray}}1=\sum_{\begin{subarray}{c}u\hskip-5.69054pt\pmod{\kappa}\\ f(u)\equiv 0\hskip-5.69054pt\pmod{\kappa}\end{subarray}}\sum_{\begin{subarray}{c}z<n\leq x\\ n\equiv u\hskip-5.69054pt\pmod{\kappa}\end{subarray}}1\leq\omega_{f}(\kappa)\left(\frac{x-z}{\kappa}+1\right),$
@@
Let $f(t)\in\mathbb{Z}[t]$ be a polynomial such that $f(t)\to\infty$ as $t\to\infty$, and $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 1$. Suppose that $f(z)>x$ is satisfied, and put h=xzh=x-z. Then we have
@@
Let $f(t)\in\mathbb{Z}[t]$ be a polynomial such that $f(t)\to\infty$ as $t\to\infty$, and $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 1$. Suppose that $f(z)>x$ is satisfied, and put h=xzh=x-z. Then inequality (2.1) holds with
@@
$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~(j=1,2)\\ P^{+}(k_{j})\leq y~(j\geq 3)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
@@
$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{i+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{i}\leq h,~\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{i}}<k_{i+1}\leq f(x)\\ P^{+}(k_{j})\leq y~(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{i+1})\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{i+1})$
@@
$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ \sqrt{y}<P^{+}(k_{1})\leq y\\ P^{+}(k_{j})\leq y~(j\geq 2)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\sum_{\begin{subarray}{c}1\leq\lambda\leq\frac{h}{k_{1}\cdots k_{m}}\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ k_{1}\cdots k_{m}\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
@@
$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m}\geq 1\\ k_{1}\cdots k_{m}\leq h\\ P^{+}(k_{j})\leq y~(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m})\sum_{\begin{subarray}{c}\frac{h}{k_{1}\cdots k_{m}}<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\omega_{f}(k_{1}\cdots k_{m}\lambda).$
@@
$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~(j=1,2)\\ P^{+}(k_{j})\leq y~(j\geq 3)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}1\leq\lambda\leq\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\sum_{\begin{subarray}{c}z<n\leq x\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\lambda\mid f(n)\\ P^{+}(f(n))\leq y\end{subarray}}1$
@@
$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq y~(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\sum_{\begin{subarray}{c}\frac{h}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}<\lambda\leq f(x)\\ P^{+}(\lambda)\leq y\end{subarray}}\Lambda(\lambda)\omega_{f}(k_{1}\cdots k_{m}\lambda).$
@@
Note that the function $\omega_{f}(k)$ is multiplicative but is not completely multiplicative in general. Hence we do not expect that $\omega_{f}(k_{1}\cdots k_{m})=\omega_{f}(k_{1})\cdots\omega_{f}(k_{m})$ holds unless $k_{1},\ldots,k_{m}$ are pairwise coprime.  [^2] It seems to be misunderstood often in [19] that the identity $\omega_{f}(k_{1}\cdots k_{m})=\omega_{f}(k_{1})\cdots\omega_{f}(k_{m})$ holds for any $k_{1},\ldots,k_{m}\in\mathbb{Z}_{\geq 1}$. One can fix the errors with the same argument presented in this paper using Lemma 4.2. Then we use the following lemma which helps us estimate multiple sums involving $\omega_{f}(k_{1}\cdots k_{m})$.
@@
Suppose that $f(t)\in\mathbb{Z}[t]$ is a product of distinct irreducible factors $f_{1}(t),\ldots,f_{g}(t)$ over $\mathbb{Z}[t]$ of degrees $d_{1},\ldots,d_{g}\geq 1$, respectively. Let $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 2$. Suppose further that $f(t)\to\infty$ as $t\to\infty$, and that $f(z)>x$. Put h=xzh=x-z. Then there exists a constant $A(f)>1$ depending only on $f(t)$ such that, for any $m\in\mathbb{Z}_{\geq 1}$ and any $i=1,\ldots,m$, we have
@@
by recalling that h=xz<xh=x-z<x. Applying (4.1) and (4.2) of Lemma 4.1, we have
@@
Suppose that $f(t)\in\mathbb{Z}[t]$ satisfies the same conditions as in Proposition 4.3. Let $x,y,z\in\mathbb{R}$ with $x>z>T_{0}(f)$ and $y\geq 2$. Put h=xzh=x-z. Then there exists a constant $A(f)>1$ depending only on $f(t)$ such that, for any $m\in\mathbb{Z}_{\geq 1}$ and any $i=1,\ldots,m$, we have
@@
$\quad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq\sqrt{y}~(j=1,2)\\ P^{+}(k_{j})\leq y~(j\geq 3)\end{subarray}}\frac{\Lambda(k_{1})\cdots\Lambda(k_{m+1})}{[k_{1},k_{2}]k_{3}\cdots k_{m+1}}\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
@@
$\qquad\times\sum_{\begin{subarray}{c}k_{1},\ldots,k_{m+1}\geq 1\\ [k_{1},k_{2}]k_{3}\cdots k_{m+1}\leq h\\ P^{+}(k_{j})\leq y~(\forall j)\end{subarray}}\Lambda(k_{1})\cdots\Lambda(k_{m+1})\omega_{f}([k_{1},k_{2}]k_{3}\cdots k_{m+1})$
@@
Then we divide the sums into those over $k_{1},\ldots,k_{m+1}\geq 1$ such that $(k_{j},\Delta_{f})=1$ if $j\in S$ and $(k_{j},\Delta_{f})>1$ otherwise, where $S$ is a subset of $\{1,\ldots,m+1\}$. Recall that $k_{1},k_{2}]k_{3}\cdots k_{m+1}=k_{1}k_{2}k_{3}\cdots k_{m+1}$ holds if $(k_{1},k_{2})=1$. Therefore we obtain
@@
$\qquad\times\frac{xg^{[u]}}{d(d-1)^{[u]-1}u^{[u]}}\left(1+O\left(\frac{1}{\log\log{x}}\right)\right)$
@@
5. Proof of Theorem 1.2 and its corollary
Let $\alpha$ be an arbitrary algebraic irrational number, and $K=\mathbb{Q}(\alpha)$ the algebraic field generated by $\alpha$ over the rationals. For any rational integer $n\geq 0$, we have
$(n+\alpha)\mathfrak{a}=\prod_{\mathfrak{p}\in J_{\alpha}}\mathfrak{p}^{u_{n}(\mathfrak{p})},$
where $J_{\alpha}$ denotes the set of all prime ideals of $K$, and $u_{n}(\mathfrak{p})$ are non-negative integers. Define $P_{\alpha}$ as the subset of $J_{\alpha}$ consisting of prime ideals $\mathfrak{p}$ such that
@@
In other words, we have $\operatorname{N}(\mathfrak{p})=p$ and $\mathfrak{p}^{2}\nmid(p)$ for a rational prime $p$ if $\mathfrak{p}\in P_{\alpha}$. Then we rewrite the factorization of $(n+\alpha)\mathfrak{a}$ as
$(n+\alpha)\mathfrak{a}=\mathfrak{b}_{n}\prod_{\mathfrak{p}\in P_{\alpha}}\mathfrak{p}^{u_{n}(\mathfrak{p})},$
where $\mathfrak{b}_{n}$ is the integral containing all prime factors of $(n+\alpha)\mathfrak{a}$ which are not in $P_{\alpha}$. By [16, Lemma 4.4], we know that there exists a constant $C(\alpha)>0$ depending at most on $\alpha$ such that
(5.1) $\operatorname{N}(\mathfrak{b}_{n})\leq C(\alpha)$
holds for any $n\in\mathbb{Z}$. Thus we have $\operatorname{N}(\mathfrak{p})\leq C(\alpha)$ for any $\mathfrak{p}\in J_{\alpha}\setminus P_{\alpha}$. Then we prove Theorem 1.2 by using the following lemmas.